How many gallons of milk containing 4% fat is mixed with milk containing 1% fat to obtain a 15 gallon mixture containing 2% fat?
Found 2 solutions by mananth, ewatrrr:
Answer by mananth(16946)

milk containing 4% fat : let this quantity be x gallons
milk containing 1% fat: this will be 15-x gallons
Mixture 15 gallon containing 2% fat?
...
0.04x+0.01(15-x)=15*0.02
0.04x+0.15-0.01x=0.3
0.03x=0.3-0.25
0.03x=0.15
x= 0.15/0.03
x= 5 gallons of 4% fat
m.ananth@hotmail.ca

Answer by ewatrrr(24785)

Hi,
Let x be the total amount of the 4% milk. The (15-x) would be the eamount of the 1% Milk
.
question states**
.04x + .01(15-x) = .02*15gal
.
simplify and solve
.04x - .01x + .01*15 = .02*15gal
.
.03x = .01*15gal
.
x = .15/.03
.
x = 5 gal
.
check your answer
.04*5 + .01*10 = .2 + .1 = .3 = .02*15